solve:
log6(5-2a)-log6(3a)=1

Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(5413)

Applying "log rules":
log6(5-2a)-log6(3a)=1
log6(5-2a)/(3a)=1
(5-2a)/(3a)=6^1
(5-2a)/(3a)=6
5-2a=18a
5 = 20a
5/20 = a
1/4 = a

Answer by stanbon(48569)

solve:
log6(5-2a)-log6(3a)=1
------------
log6[(5-2a)/(3a)] = 1
(5-2a)/3a = 6^1
5-2a = 18a
20a = 5
a = 1/4
-----------------
Cheers,
Stan H.